OpenStudy (anonymous):

What is the frequency of light with a wavelength of 7.00x10^-7 m

OpenStudy (anonymous):

The velocity of a wave is equal to it's wavelength multiplied by its frequency, in the case of electromagnetic waves, we know they all travel at c, the speed of light, so if we know either the wavelength of frequency we can solve for the other.

OpenStudy (anonymous):

This is what I Put for my answer: (3x10^10)(7.00X10^-7)=21x10^3

OpenStudy (anonymous):

This isn't quite right. \[c = \lambda f = 3 * 10^{8} = 7.00 *10^{-7} f\] just solve for f!

OpenStudy (anonymous):

I'm sorry I am still confused. Do I divide 7.00*10^-7 from both sides?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

Thank you :)